A cubical block of side 30 cm is moving with velocity  2 m/s on a smooth surface . The surface has a bump. The angular velocity of block immediately after it hits the bump is?

Grade:11

1 Answers

Arun
6 years ago
Dear Mohan
Using conservation of angular momentum
m*v* (0.3/2) = (2/3)* m*(0.3)²*w
w = (3*2)/(4*0.3)
w = 5 rad/sec
Regards
Arun (askIITians forum expert)